A new machine has just been installed to cut and rough-shape large slugs. The slugs are then transferred to a precision grinder. One of the critical measurements is the outside diameter. The quality control inspector randomly selected five slugs each half-hour, measured the outside diameter, and recorded the results. The measurements (in millimeters) for the period 8:00 a.m. to 10:30 a.m. follow.

Outside Diameter (millimeters)

Time

1

2

3

4

5

8:00

87.1

87.4

88.1

87.0

87.0

8:30

86.9

88.5

87.6

87.5

87.4

9:00

87.6

88.4

86.8

87.8

88.2

9:30

86.0

88.0

87.2

87.6

87.1

10:00

87.2

86.9

87.3

87.1

86.8

10:30

88.1

86.2

87.3

87.3

87.9

  1. Determine the control limits for the mean and the range.(Leave no cells blank - be certain to enter "0" wherever required. Round your answers to 2 decimal places.)

Mean of the sample means

=

LCL

=

UCL

=

Mean of the sample ranges

=

LCL

=

UCL

=

  1. c-1. Are any means out of control?

Yes

No

  1. c-2. Are any ranges out of control?

Yes

No
Answer #1

Solution:

. 31 u , 5 | Mean 874 1881 | 87 | 87 | 87.32 | Range Tos 8:00 87. 1 86.9 87.58 87.6 87.76 87-18 86 10:00 87.2 10:36 88.1 |86.UCL = šta R 2 87.390 +(0.577) *(1.45). UCL * 88.220. LOL LEX-A R = 87.390 - 20.571)*(1-45) LoL – 86 * chart UCL = 88-228 - 828 chall 3.065 ~ 3.07 RCL = Du R = 2:114*1:45 * LCL = DJ R 0 * 1.45 =0. R =1.45 Rchalt UCL = 3.000 R=1.45 - LOL=0

c-1) Solution: No

c-2) Solution: No